OpenStudy (anonymous):

Find the term in x^k for the indicated value of k in the expansion of (x+x^2)^3 k=2

OpenStudy (anonymous):

\[(x+x^2)^3\] find the term in \[x^2\]

OpenStudy (anonymous):

When i calculated using the the General term, I got does not exist since i got r=-1

OpenStudy (anonymous):

\[GT=13Cr(6x)^{13-r}(x^2)^{r}\]

OpenStudy (anonymous):

My bad that should be 3 no 13

OpenStudy (anonymous):

oh woops im reading the wrong answer for the question lol

OpenStudy (aum):

I would factor out the x which will become x^3 outside. Will there ever be an x^2 term?

OpenStudy (aum):

(x+x^2)^3 = {x(1+x)}^3 = x^3 * (1+x)^3

OpenStudy (anonymous):

\[GT=3Cr(x)^{3-r}(x^2)^r\] \[=3Cr(x)^{3+r}\]<--when simiplified

OpenStudy (anonymous):

using the exponents law i did 3+r=2 and got r=-1 the term DNE, can someone confirm that?

OpenStudy (aum):

Read my reply above.

OpenStudy (anonymous):

right so it does not exist right

OpenStudy (aum):

x^3 * (1+x)^3 The lowest exponent after expansion will be 3. There will be no x^2 term. It does not exist.

OpenStudy (anonymous):

alright thanks

OpenStudy (aum):

yw.